# cammarl: Conformal Action Modeling in Multi Agent Reinforcement Learning

### v1

@@ Abstract
Before taking actions in an environment with more than one intelligent agent, an autonomous agent may benefit from reasoning about the other agents and utilizing a notion of a guarantee or confidence about the behavior of the system. In this article, we propose a novel multi-agent reinforcement learning (MARL) algorithm cammarl, which involves modeling the actions of other agents in different situations in the form of confident sets, i.e., sets containing their true actions with a high probability. We then use these estimates to inform an agent’s decision-making. For estimating such sets, we use the concept of conformal predictions, by means of which, we not only obtain an estimate of the most probable outcome but get to quantify the operable uncertainty as well. For instance, we can predict a set that provably covers the true predictions with high probabilities (e.g., 95%). Through several experiments in two fully cooperative multi-agent tasks, we show that cammarl elevates the capabilities of an autonomous agent in MARL by modeling conformal prediction sets over the behavior of other agents in the environment and utilizing such estimates to enhance its policy learning. All developed codes can be found here: https://github.com/Nikunj-Gupta/conformal-agent-modelling.
@@ 1 Introduction
Developing systems of autonomous agents capable of effective multi-agent interactions can be very useful in modern cooperative artificial intelligence (AI). For instance, service robots, surveillance agents, and many more similar applications require profound collaboration among agents (and with humans), without prior coordination. Now, to enable complex, constructive behaviors to emerge from unsupervised interactions among agents, an essential skill for an agent to have is the ability to reason about other agents in the environment. There has been considerable research addressing this problem of agent or opponent modeling [2]. Generally, it involves constructing models of other agents that learn useful attributes to inform its own decision-making (such as the future actions of the other agents, or their current goals and plans) from current or past interaction history (such as the previous actions taken by other agents in different situations).
We are interested in the particular aspect of an interactive, autonomous agent which involves learning an additional, independent model to make predictions about the actions of the other agents in the environment, supplemental to its reinforcement learning-based policy to make decisions related to its downstream task. An autonomous agent can then incorporate those estimates to inform its decision-making and optimize its interaction with the other agents. While there exist several methods for developing such models for other agents  [2], there is currently no method or theory to the best of our knowledge which would allow an agent to consider the correctness or confidence of the predictions of the learned model.
Figure 1: Our proposed methodology of informing an autonomous agent’s decision-making by means of conformal predictions of action sets of other agents in the environment. Two agents ($\mathcal{N}_{self}$, $\mathcal{N}_{other}$) receive their own partial observations from the environment ($o_{self}$, $o_{other}$) and take their actions ($a_{self}$, $a_{other}$). An independent conformal action prediction model $\mathcal{C}$ learns to output a conformal action set, $\{a^{\prime}_{other}\}$, corresponding to $\mathcal{N}_{other}$ which are then used as additional inputs for training by $\mathcal{N}_{self}$ to inform its policy and perform its action $a_{self}$.
@@
Conformal predictions or inference is a fitting method for generating statistically accurate uncertainty sets for the predictions from machine learning classifiers. It is steadily gaining popularity owing to its explicit and non-asymptotic guarantees over the produced sets [5]. In other words, we can obtain conformal sets that provably contain the true predictions with high probabilities, such as 95%, chosen in advance. This can be very useful and successful in high-risk learning settings, especially in decision-making in medical applications from diagnostic information, for instance, which demand quantifying uncertainties to avoid insufferable model failures. What if we only prefer to use the predictions when the model is confident? For example, doctors may only consider a predicted medical diagnosis when the model is at least 95% accurate, or may want to use the predicted set with high credence to consider ruling out relevant possibilities. So, in this article, we aim to enhance the capabilities of an agent in a multi-agent reinforcement learning (MARL) setting by modeling and using conformal prediction sets (or the latent representations learned in the process [^1] More details in Appendix A.) over the behavior of an autonomous system. In particular, we model other agents’ actions in the form of confident sets, i.e., sets that contain other agents’ true actions with a high probability. We hypothesize that these estimated conformal sets would inform our learning agent’s decision-making and elevate its performance in MARL. Figure 1 shows the high-level idea of our proposed model for learning agents in any given environment.
@@ 2 Related Works
Decision-making without reasoning about other agents in the environment can be very challenging, for instance, due to weak or no theoretical guarantees, non-stationarity (single agent’s perspective), and inefficient coordination for a considerable coherent joint behavior [25]. Modeling other agents in an environment is not new and has been studied in the past [2, 3]. However, our proposal of predicting conformal sets of actions of the other agents in the environment (with high probability) is novel and has not been attempted to the best of our knowledge.
@@
Model-based reinforcement learning (MBRL) has certainly shown its advantages in data efficiency, generalization, exploration, counterfactual reasoning, and performance in many tasks and domains [14, 15, 18, 28, 31, 38] in single-agent RL, and now, it has also started to attract attention in MARL [44]. However, most of the current works in model-based MARL do not yet focus on teammate or opponent modeling. Some recent works [36, 46] incorporated dynamics modeling and a prediction module to estimate the actions of other agents within the construction of the environment model. However, these prediction models were trained without accessing the true trajectories from the other agents which can be problematic in several use cases.
@@
A widely popular technique to reason about other agents in the environment is to learn representations of different properties of other agents. For instance, learning to reconstruct the actions of other agents from their partial observations [16, 26, 32, 1], modeling an agent or its policy using encoder-decoder-based architectures [12, 48], learning latent representations from local information with or without utilizing the modeled agent’s trajectories [34, 45] or modeling the forward dynamics of the system through relational reasoning using graph neural networks [42]. Theory-of-Mind Network or TomNet learned embeddings corresponding to other agents in the environment for meta-learning [39]. Some works also constructed I-POMDPs to utilize recursive reasoning [2] assuming unrestricted knowledge of the observation models of other agents. Nevertheless, cammarl involves no form of reconstruction of other agent’s policy or rewards, or state models. Any of these techniques can be used with cammarl  which, however, is not the objective of this work. Also, unlike cammarl, many of these aforementioned techniques evaluate in fully-observable environments or rely upon direct access to other agents’ experience trajectories even during execution. This can be infeasible in various settings.
@@
Numerous deep MARL research works that focus on partial observability in fully cooperative settings indirectly involve reasoning about the intentions of teammates or opponents in an environment [11]. For instance, many works allow agents to communicate, enabling them to indirectly reason about the others’ intentions [20, 9, 41, 8, 30, 13, 47]. On the other hand, some studied the emergence of cooperative and competitive behaviors among agents in varying environmental factors, for instance, task types or reward structures [22]. Recent work in hierarchical reinforcement learning also attempts to develop a hierarchical model to enable agents to strategically decide whether to cooperate or compete with others in the environment and then execute respective planning programs [19]. However, none of these works study the improvement in an autonomous agent’s decision-making via directly modeling the other agents in the environment or predicting their actions or current or future intentions.
@@
Research in the field of IRL also relates to our work because we share the key motive of inferring other agents’ intentions and then use it to learn a policy that maximizes the utility of our learning agent [6]. However, IRL addresses this by deducing the reward functions of other agents based on their behavior, assuming it to be nearly optimal. On the other hand, in cammarl  we directly model the other agent’s actions based on their observations and use these estimates to indirectly infer their goal in an online manner.
@@
Estimating well-grounded uncertainty in predictions is a difficult and unsolved problem and there have been numerous approaches for approximating it in research in supervised learning [10]. Recent works in conformal predictions [4, 21, 17, 35, 7, 27] have now significantly improved upon some of the early research [43, 37, 33], for instance in terms of predicted set sizes, improved efficiency, and providing formal guarantees. For this article, we adapt the core ideas from Regularized Adaptive Prediction Sets (RAPS) [4] to our setting owing to its demonstrated improved performance evaluation on classification benchmarks in supervised learning [4].
@@ 3.1 Mathematical Model
Formally, we consider two agents in the environment — learning agent denoted by self and the other agent denoted by other. The partially observable Markov game [23] for our setting can then be defined using the following tuple [^2] A tabular version can be found in Appendix C.:
$\langle\mathcal{N}_{i},\mathcal{S},\mathcal{A}_{i},\mathcal{O}_{i},\mathcal{T},\mathcal{C},\pi_{\theta_{i}},r_{i}\rangle_{i\in\{self,other\}}$
With the set $\mathcal{S}$ describing the possible true states (or full observations) of the environment, two agents, $\mathcal{N}_{self}$ and $\mathcal{N}_{other}$, observe the environment locally using their sets of observations $\mathcal{O}_{self}$ and $\mathcal{O}_{other}$ respectively, and act using their set of actions, $\mathcal{A}_{self}$ and $\mathcal{A}_{other}$. Each agent i can select an action $a_{i}\in\mathcal{A}_{i}$ using their policy $\pi_{\theta_{i}}$, and their joint action $\textbf{a}\in\mathcal{A}_{self}\times\mathcal{A}_{other}$ then imposes a transition to the next state in the environment according to the state transition function $\mathcal{T}$, defined as a probability distribution on the subsequent state based on current state and actions, $\mathcal{T}$ : $\mathcal{S}\times\mathcal{A}_{self}\times\mathcal{A}_{other}\times\mathcal{S}\rightarrow[0,1]$. The agents use their individual reward function $r_{i}(s,a)$ : $\mathcal{O}_{i}\times\mathcal{A}_{i}\rightarrow\mathbb{R}$. Both agents aim to maximize their own total expected rewards $R_{i}=\sum_{t=0}^{T}\gamma^{t}r_{i}^{t}$ where $\gamma\in[0,1)$ as the discount factor and T is the time horizon.
In cammarl, at each time step t, we also use a conformal prediction model defined as a set-valued function $\mathcal{C}:\mathbb{R}^{d}\rightarrow 2^{\mathcal{A}_{other}}$
$\mathcal{C}(o^{t}_{other})\rightarrow\{A^{i}_{other}\}$
which outputs a conformal action predictive set $\{A^{t}_{other}\}$ for each input of $\mathcal{N}_{other}$’s local observation $o^{t}_{other}\in\mathcal{O}_{other}$ at the time.
@@ 3.2 Conformal Action Modeling
1:$N_{self}$, $N_{other}$leftarrow Initialize Actor-Critic networks for each — $\mathcal{N}_{self}$ and $\mathcal{N}_{other}$
2:$conformalModel$leftarrow Initialize the conformal model to predict $\mathcal{N}_{other}$’s conformal action sets
3:$b_{self}$, $b_{other}$, $b_{conformal}$leftarrow experience replay memory buffers for $\mathcal{N}_{self}$, $\mathcal{N}_{other}$ $conformalModel$
4:for $episode=1,2,\ldots$  do
5:     Fetch observations $o_{self}$, $o_{other}$ from environment
6:     for $timesteps=1,2,\ldots,T$  do
7:         $conformalActions\leftarrow conformalModel(o_{other})$triangleright Predict $\mathcal{N}_{other}$’s conformal set of actions
8:         $o_{self}$leftarrow $o_{self}$ + $conformalActions$triangleright Concatenate conformal actions to $o_{self}$
9:         Run policies $\pi_{\theta_{self}}$ and $\pi_{\theta_{other}}$ in the environment
10:         Collect trajectories of $\mathcal{N}_{self}$ and $\mathcal{N}_{other}$ in $b_{self}$ and $b_{other}$
11:         Collect $\mathcal{N}_{other}$’s state-action mappings in $b_{conformal}$
12:         if  update interval reached  then
13:              Train $conformalModel$ using $b_{conformal}$
14:              Train $N_{self}$ on $b_{self}$ using PPO
15:              Train $N_{other}$ on $b_{other}$ using PPO
16:         end if
17:     end for
18:end for
@@
Now we formally describe our proposed algorithm — Conformal Action Modeling-based Multi-Agent Reinforcement Learning or cammarl. Our objective is to inform an $\mathcal{N}_{self}$’s decision-making by modeling $\mathcal{N}_{other}$’s actions in the environment as conformal prediction sets which contain the true actions with a high probability (for example, 95%). More specifically, $\mathcal{N}_{self}$ uses a separate conformal action prediction model to obtain sets of $\mathcal{N}_{other}$’s actions at each timestep that contains latter’s true action in the environment at a given time step with high prespecified probabilities.
@@
Algorithm 1 describes the complete workflow of training of agents in cammarl. We begin by initializing the actor-critic networks for both the agents in the environment, the conformal model, and the memory buffers for each of these (lines 1-3). Now, at the beginning of each episode in the environment, both agents receive their own partial observations (line 5). Next, the conformal model predicts the actions of $\mathcal{N}_{other}$ in the form of a set, which is then provided as an additional input to $\mathcal{N}_{self}$ (lines 7-8), whereas $\mathcal{N}_{other}$ has access to only its own partial observation, $o_{other}$. Both agents now take actions in the environment and continue collecting their experiences (lines 9-11). The agents and the conformal model periodically train using their respective experience memory (lines 12-16).
@@
Figure 2 shows a detailed illustration of our conformal action modeling that materializes internally at each time step. The conformal predictor $\mathcal{C}$ collects $\mathcal{N}_{other}$’s state-action pairs and periodically learns and updates a neural network classifier, $f(\cdot):\mathbb{R}^{d}\rightarrow\mathbb{R}^{|\mathcal{A}_{other}|}$ (where d is the number of dimensions in $\mathcal{N}_{other}$’s local observation and $|\mathcal{A}_{other}|$ is the number of possible discrete actions available for $\mathcal{N}_{other}$), to predict action from a given state. Then, we adapt RAPS conformal calibration [4] to our setting. Considering $\textbf{o}\in O_{other}$ as feature vectors, we use the updated f to compute action probabilities $\hat{\pi}_{o}\in\mathbbm{R}^{|\mathcal{A}_{other}|}$. The probabilities are then ordered from most probable to least probable followed by the estimation of the predictive action set for the given feature inputs. To promote small predictive sets we also add a regularization term as also proposed in RAPS. Formally, for a feature vector o and the corresponding possible prediction a, to estimate a set which includes all the actions that will be included before a, let us define the total probability mass of the set of actions that are more probable than a:
@@
where $(x)^{+}$ denotes the positive portion of x, and $\lambda,k_{reg}\geq 0$ are regularization hyperparameters to incentivize small set sizes. Here, $u\sim uniform$ [0, 1] (used to allow for randomized procedures) and the tuning parameter $\tau$ (the cumulative sum of the classifier scores after sorting and penalization) to control the size of the sets are identical to as used in RAPS for supervised tasks [4].
@@
4 Environments
In this section, we discuss the cooperative tasks with two agents used in this study (Figure 3). We note here that though we work in fully cooperative settings in this article, cammarl as an idea can be generalized to competitive or mixed settings too.
(a) Cooperative Navigation: Agents (blue) learn to cover the two landmarks (black) avoiding collisions.
(b) Level-based foraging: Agents must collect food and learn to cooperate using sparse rewards.
Figure 3: Multi-agent cooperative environments used in this study: (a) OpenAI MPE’s cooperative navigation with 2 agents (N=2) and 2 landmarks (L=2) and (b) a 12 $\times$ 12 level-based foraging grid-world with 2 cooperative players and 4 food locations.
Cooperative Navigation (CN) [29, 24].
where $|\mathcal{N}|$ is the number of agents and L are the landmarks in the environment. Here, c is the number of collisions in an episode and the agents are penalized with -1 for each time two agents collide. Figure 2(a) shows an illustration of this environment.
Level-based foraging (LBF) [1].
In this environment, $\mathcal{N}_{self}$ and $\mathcal{N}_{other}$ are part of a 12$\times$12 grid world which contains four randomly scattered food locations, each assigned a level. The agents also have a level of their own. They attempt to collect food which is successful only if the sum of the levels of the agents involved in loading is greater than or equal to the level of the food. This is a challenging environment, requiring agents to learn to trade off between collecting food for their own and cooperating with the other agent to acquire higher team rewards. Moreover, this environment has sparse rewards making it difficult to learn and operate independently in the environment. In particular, each agent is rewarded equal to the level of the food it managed to collect, divided by its level (its contribution). Figure 2(b) shows an illustration of this environment.
5 Experiments and Results
(a) CN
(b) LBF
Figure 4: Comparison of agent performances (in terms of reward accumulation) in CN and LBF in different settings with varying pieces of information available to $\mathcal{N}_{self}$ during training. cammarl’s performance is very close to the upper bound, GIAM, and is considerably better than the other extreme, NOAM. It also outperforms the other defined benchmarks (TAAM, TOAM, & EAP) in both tasks, along with the benefit of uncertainty quantification of its estimates. Interestingly, in CN, cammarl can be seen to learn arguably faster, but all methods converge to similar results, whereas in LBF, it actually seems to converge to a better policy. The curves are averaged over five independent trials and smoothed using a moving window average (100 points) for readability.

### v2

@@ Abstract
Before taking actions in an environment with more than one intelligent agent, an autonomous agent may benefit from reasoning about the other agents and utilizing a notion of a guarantee or confidence about the behavior of the system. In this article, we propose a novel multi-agent reinforcement learning (MARL) algorithm cammarl, which involves modeling the actions of other agents in different situations in the form of confident sets, i.e., sets containing their true actions with a high probability. We then use these estimates to inform an agent’s decision-making. For estimating such sets, we use the concept of conformal predictions, by means of which, we not only obtain an estimate of the most probable outcome but get to quantify the operable uncertainty as well. For instance, we can predict a set that provably covers the true predictions with high probabilities (e.g., 95%). Through several experiments in two fully cooperative multi-agent tasks, we show that cammarl elevates the capabilities of an autonomous agent in MARL by modeling conformal prediction sets over the behavior of other agents in the environment and utilizing such estimates to enhance its policy learning. Implementation can be found here: https://github.com/Nikunj-Gupta/conformal-agent-modelling.
Machine Learning, ICML
@@ 1 Introduction
Developing systems of autonomous agents capable of effective multi-agent interactions can be very useful in modern cooperative artificial intelligence (AI). For instance, service robots, surveillance agents, and many more similar applications require profound collaboration among agents (and with humans), without prior coordination. Now, to enable complex, constructive behaviors to emerge from unsupervised interactions among agents, an essential skill for an agent to have is the ability to reason about other agents in the environment. There has been considerable research addressing this problem of agent or opponent modeling (Albrecht & Stone, 2018). Generally, it involves constructing models of other agents that learn useful attributes to inform its own decision-making (such as the future actions of the other agents, or their current goals and plans) from current or past interaction history (such as the previous actions taken by other agents in different situations).
Figure 1: Our proposed methodology of informing an autonomous agent’s decision-making by means of conformal predictions of action sets of other agents in the environment illustrated with two agents for simplicity. Two agents ($\mathcal{N}_{self}$, $\mathcal{N}_{other}$) receive their own partial observations from the environment ($o_{self}$, $o_{other}$) and take their actions ($a_{self}$, $a_{other}$). An independent conformal action prediction model $\mathcal{C}$ learns to output a conformal action set, $\{a^{\prime}_{other}\}$, corresponding to $\mathcal{N}_{other}$ which are then used as additional inputs for training by $\mathcal{N}_{self}$ to inform its policy and perform its action $a_{self}$.
We are interested in the particular aspect of an interactive, autonomous agent that involves learning an additional, independent model to make predictions about the actions of the other agents in the environment, supplemental to its reinforcement learning-based policy to make decisions related to its downstream task. An autonomous agent can then incorporate those estimates to inform its decision-making and optimize its interaction with the other agents. While there exist several methods for developing such models for other agents (Albrecht & Stone, 2018), there is currently no method or theory to the best of our knowledge that would allow an agent to consider the correctness or confidence of the predictions of the learned model.
@@
Conformal predictions or inference is a fitting method for generating statistically accurate uncertainty sets for the predictions from machine learning classifiers. It is steadily gaining popularity owing to its explicit and non-asymptotic guarantees over the produced sets (Angelopoulos & Bates, 2021). In other words, we can obtain conformal sets that provably contain the true predictions with high probabilities, such as 95%, chosen in advance. This can be very useful and successful in high-risk learning settings, especially in decision-making in medical applications from diagnostic information, for instance, which demand quantifying uncertainties to avoid insufferable model failures. What if we only prefer to use the predictions when the model is confident? For example, doctors may only consider a predicted medical diagnosis when the model is at least 95% accurate, or may want to use the predicted set with high credence to consider ruling out relevant possibilities. So, in this article, we aim to enhance the capabilities of an agent in a multi-agent reinforcement learning (MARL) setting by modeling and using conformal prediction sets (or the latent representations learned in the process [^1] More details in Appendix B.) over the behavior of an autonomous system. In particular, we model other agents’ actions in the form of confident sets, i.e., sets that contain other agents’ true actions with a high probability. We hypothesize that these estimated conformal sets would inform our learning agent’s decision-making and elevate its performance in MARL. Figure 1 shows the high-level idea of our proposed model for learning agents in any given environment.
@@ 2 Related Works
Decision-making without reasoning about other agents in the environment can be very challenging, for instance, due to weak or no theoretical guarantees, non-stationarity (single agent’s perspective), and inefficient coordination for a considerable coherent joint behavior (Matignon et al., 2012). Modeling other agents in an environment is not new and has been studied in the past (Albrecht & Stone, 2018; Albrecht et al., 2020). However, our proposal of predicting conformal sets of actions of the other agents in the environment (with high probability) is novel and has not been attempted to the best of our knowledge.
@@
Model-based reinforcement learning (MBRL) has certainly shown its advantages in data efficiency, generalization, exploration, counterfactual reasoning, and performance in many tasks and domains (Hafner et al., 2020, 2021; Jain et al., 2022; Moerland et al., 2023; Pal & Leon, 2020; Polydoros & Nalpantidis, 2017) in single-agent RL, and now, it has also started to attract attention in MARL (Wang et al., 2022). However, most of the current works in model-based MARL do not yet focus on teammate or opponent modeling. Some recent works (Park et al., 2019b; Zhang et al., 2021) incorporated dynamics modeling and a prediction module to estimate the actions of other agents within the construction of the environment model. However, these prediction models were trained without accessing the true trajectories from the other agents which can be problematic in several use cases.
@@
A widely popular technique to reason about other agents in the environment is to learn representations of different properties of other agents. For instance, learning to reconstruct the actions of other agents from their partial observations (He et al., 2016; Mealing & Shapiro, 2015; Panella & Gmytrasiewicz, 2017; Albrecht & Ramamoorthy, 2015), modeling an agent or its policy using encoder-decoder-based architectures (Grover et al., 2018; Zintgraf et al., 2021), learning latent representations from local information with or without utilizing the modeled agent’s trajectories (Papoudakis et al., 2021; Xie et al., 2021) or modeling the forward dynamics of the system through relational reasoning using graph neural networks (Tacchetti et al., 2018). Theory-of-Mind Network or TomNet learned embeddings corresponding to other agents in the environment for meta-learning (Rabinowitz et al., 2018). Some works also constructed I-POMDPs to utilize recursive reasoning (Albrecht & Stone, 2018) assuming unrestricted knowledge of the observation models of other agents. Nevertheless, cammarl involves no form of reconstruction of other agent’s policy or rewards, or state models. Any of these techniques can be used with cammarl  which, however, is not the objective of this work. Also, unlike cammarl, many of these aforementioned techniques evaluate in fully-observable environments or rely upon direct access to other agents’ experience trajectories even during execution. This can be infeasible in various settings.
@@
Numerous deep MARL research works that focus on partial observability in fully cooperative settings indirectly involve reasoning about the intentions of teammates or opponents in an environment (Gronauer & Diepold, 2022). For instance, many works allow agents to communicate, enabling them to indirectly reason about the others’ intentions (Lazaridou et al., 2016; Foerster et al., 2016; Sukhbaatar et al., 2016; Das et al., 2017; Mordatch & Abbeel, 2018; Gupta et al., 2021; Zhu et al., 2022). On the other hand, some studied the emergence of cooperative and competitive behaviors among agents in varying environmental factors, for instance, task types or reward structures (Leibo et al., 2017). Recent work in hierarchical reinforcement learning also attempts to develop a hierarchical model to enable agents to strategically decide whether to cooperate or compete with others in the environment and then execute respective planning programs (Kleiman-Weiner et al., 2016). However, none of these works study the improvement in an autonomous agent’s decision-making via directly modeling the other agents in the environment or predicting their actions or current or future intentions.
@@
Research in the field of IRL also relates to our work because we share the key motive of inferring other agents’ intentions and then use it to learn a policy that maximizes the utility of our learning agent (Arora & Doshi, 2021). However, IRL addresses this by deducing the reward functions of other agents based on their behavior, assuming it to be nearly optimal. On the other hand, in cammarl  we directly model the other agent’s actions based on their observations and use these estimates to indirectly infer their goal in an online manner.
@@
Estimating well-grounded uncertainty in predictions is a difficult and unsolved problem and there have been numerous approaches for approximating it in research in supervised learning (Gawlikowski et al., 2021). Recent works in conformal predictions (Angelopoulos et al., 2020; Lei et al., 2018; Hechtlinger et al., 2018; Park et al., 2019a; Cauchois et al., 2020; Messoudi et al., 2020) have now significantly improved upon some of the early research (Vovk et al., 2005; Platt et al., 1999; Papadopoulos et al., 2002), for instance in terms of predicted set sizes, improved efficiency, and providing formal guarantees. For this article, we adapt the core ideas from Regularized Adaptive Prediction Sets (RAPS) (Angelopoulos et al., 2020) to our setting owing to its demonstrated improved performance evaluation on classification benchmarks in supervised learning (Angelopoulos et al., 2020).
@@ 3.1 Mathematical Model
Formally, we consider two agents in the environment — learning agent denoted by self and the other agent denoted by other. The partially observable Markov game (Littman, 1994) for our setting can then be defined using the following tuple [^2] A tabular version can be found in Section C.1:
$\langle\mathcal{N}_{i},\mathcal{S},\mathcal{A}_{i},\mathcal{O}_{i},\mathcal{T},\mathcal{C},\pi_{\theta_{i}},r_{i}\rangle_{i\in\{self,other_{1}\dots other_{K-1}\}}$
With the set $\mathcal{S}$ describing the possible true states (or full observations) of the environment, $K$ agents, $\mathcal{N}_{self}$ and $(K-1)$ $\mathcal{N}_{other_{j}}$ ($j\in[1,K]$), observe the environment locally using their sets of observations $\mathcal{O}_{self}$ and $(K-1)$ $\mathcal{O}_{other_{j}}$ respectively, and act using their set of actions, $\mathcal{A}_{self}$ and $(K-1)$ $\mathcal{A}_{other_{j}}$. Each agent i can select an action $a_{i}\in\mathcal{A}_{i}$ using their policy $\pi_{\theta_{i}}$, and their joint action $\textbf{a}\in\mathcal{A}_{self}\times\mathcal{A}_{other_{1}}\times\dots\mathcal{A}_{other_{K}}$ then imposes a transition to the next state in the environment according to the state transition function $\mathcal{T}$, defined as a probability distribution on the subsequent state based on current state and actions, $\mathcal{T}$: $\mathcal{S}\times\mathcal{A}_{self}\times\mathcal{A}_{other_{1}}\times\dots\mathcal{A}_{other_{K}}\times\mathcal{S}\rightarrow[0,1]$. The agents use their individual reward function $r_{i}(s,a)$ : $\mathcal{O}_{i}\times\mathcal{A}_{i}\rightarrow\mathbb{R}$. Both agents aim to maximize their own total expected rewards $R_{i}=\sum_{t=0}^{T}\gamma^{t}r_{i}^{t}$ where $\gamma\in[0,1)$ as the discount factor and T is the time horizon.
In cammarl, at each time step t, we also use a conformal prediction model for the $j$-th agent, defined as a set-valued function, $\mathcal{C}:\mathbb{R}^{d}\rightarrow 2^{|\mathcal{A}_{other_{j}}|}$
$\mathcal{C}(o^{t}_{other_{j}})\rightarrow\{A^{i}_{other_{j}}\}$
which outputs a conformal action predictive set $\{A^{t}_{other_{j}}\}$ for each input of $\mathcal{N}_{other_{j}}$’s local observation $o^{t}_{other_{j}}\in\mathcal{O}_{other_{j}}$ at the time.
@@ 3.2 Conformal Action Modeling
[…]
$\texttt{episode}=1,2,\ldots$
Fetch observations $o_{self}$, $o_{other_{1}}\dots o_{other_{K}}$ from environment
$\texttt{timesteps}=1,2,\ldots,T$
[…]
Collect trajectories of $\mathcal{N}_{self}$ and $\mathcal{N}_{other_{1}}\dots\mathcal{N}_{other_{K}}$
[…]
@@
Now we formally describe our proposed algorithm — Conformal Action Modeling-based Multi-Agent Reinforcement Learning or cammarl. Our objective is to inform an $\mathcal{N}_{self}$’s decision-making by modeling the other agent’s actions in the environment as conformal prediction sets that contain the true actions with a high probability (for example, 95%). More specifically, $\mathcal{N}_{self}$ uses a separate conformal action prediction model to obtain sets of $\mathcal{N}_{other_{j}}$’s actions at each timestep that contains latter’s true action in the environment at a given time step with high prespecified probabilities.
@@
Algorithm 1 describes the complete workflow of training of agents in cammarl. We begin by initializing the actor-critic networks for both the agents in the environment, the conformal model, and the memory buffers for each of these. Now, at the beginning of each episode in the environment, both agents receive their own partial observations (line 1). Next, the conformal model predicts the actions of all the $\mathcal{N}_{other_{j}}$’s in the form of a set, which is then provided as an additional input to $\mathcal{N}_{self}$ (lines 1-1), whereas $\mathcal{N}_{other_{j}}$ has access to only its own partial observation, $o_{other_{j}}$. Both agents now take actions in the environment and continue collecting their experiences (lines 1-1). The agents and the conformal model periodically train using their respective experience memory (lines 1-1).
@@
Figure 2 shows a detailed illustration of our conformal action modeling that materializes internally at each time step. We use only one other agent $\mathcal{N}_{other}$ for simplicity. The conformal predictor $\mathcal{C}$ collects $\mathcal{N}_{other}$’s state-action pairs and periodically learns and updates a neural network classifier, $f(\cdot):\mathbb{R}^{d}\rightarrow\mathbb{R}^{|\mathcal{A}_{other}|}$ (where d is the number of dimensions in $\mathcal{N}_{other}$’s local observation and $|\mathcal{A}_{other}|$ is the number of possible discrete actions available for $\mathcal{N}_{other}$), to predict action from a given state. Then, we adapt RAPS conformal calibration (Angelopoulos et al., 2020) to our setting. Considering $\textbf{o}\in O_{other}$ as feature vectors, we use the updated f to compute action probabilities $\hat{\pi}_{o}\in\mathbbm{R}^{|\mathcal{A}_{other}|}$. The probabilities are then ordered from most probable to least probable followed by the estimation of the predictive action set for the given feature inputs. To promote small predictive sets we also add a regularization term as also proposed in RAPS. Formally, for a feature vector o and the corresponding possible prediction a, to estimate a set which includes all the actions that will be included before a, let us define the total probability mass of the set of actions that are more probable than a:
@@
where $(x)^{+}$ denotes the positive portion of x, and $\lambda,k_{reg}\geq 0$ are regularization hyperparameters to incentivize small set sizes. Here, $u\sim uniform$ [0, 1] (used to allow for randomized procedures) and the tuning parameter $\tau$ (the cumulative sum of the classifier scores after sorting and penalization) to control the size of the sets are identical to as used in RAPS for supervised tasks (Angelopoulos et al., 2020).
@@
4 Experiments
In this section, we discuss the cooperative tasks with multiple agents used in this study. We note here that though we work in fully cooperative settings in this article, cammarl as an idea can be generalized to competitive or mixed settings too (more on this in Appendix).
4.1 Environments
We focus on four cooperative multi-agent environments: Cooperative Navigation The agents learn to visit the two landmarks avoiding collisions. In Level-based Foraging, agents collect foods based on their levels. In Pressure Plate, the agents traverse a grid world and must cooperate by standing on a plate to keep a gate open while the other agents collect the main reward. Finally, we also run cammarlin Google Football, where 3 agents try to score a goal against a defender and a keeper in a game of football. Further details on these environments are in Section A.
4.2 Results
Figure 3: Comparison of agent performances (in terms of reward accumulation) in CN (a) and LBF (b) in different settings with varying pieces of information available to $\mathcal{N}_{self}$ during training. cammarl’s performance is very close to the upper bound, GIAM, and is considerably better than the other extreme, NOAM. It also outperforms the other defined benchmarks (TAAM, TOAM, & EAP) in both tasks, along with the benefit of uncertainty quantification of its estimates. Interestingly, in CN, cammarl can be seen to learn arguably faster, but all methods converge to similar results, whereas in LBF, it actually seems to converge to a better policy. The curves are averaged over five independent trials and smoothed using a moving window average (100 points) for readability.